OpenStudy (anonymous):

Solve the system y = x + 2 and y = –x – 3 by using graph paper or graphing technology.

OpenStudy (anonymous):

lol. OpenStudy doesn't strike me as graph paper ;)

OpenStudy (anonymous):

is OpenStudy.com technology?  x = -5/2 y = -1/2 Answer: (-5/2, -1/2)